# A Guide to Tracking Phylogenies in Parallel and Distributed Agent-based Evolution Models

### v1

@@ Introduction
Recent advances in computing hardware harbor significant untapped potential to unleash transformative, orders-of-magnitude growth in the scale and sophistication of agent-based evolution modeling and application-oriented evolutionary computation. For instance, emerging AI/ML accelerator hardware platforms (e.g., Cerebras Wafer-Scale Engine, Graphcore Intelligence Processing Unit) currently afford up to hundreds of thousands of processing cores within a single device (lauterbach2021path; jia2019dissecting). While these hardware platforms bear limitations characteristic of highly distributed, many-processor computation, their architecture is well-suited for work with agent-based models (ABM); physical constraints in the layout of processor cores mirror the locally structured interactions typical in ABM. Nevertheless, significant challenges must be solved to effectively harness highly distributed, many-processor computation for ABM workloads.
To advance on this front, we propose a fundamental re-frame of simulation that shifts from a paradigm of “complete,” deterministic observability of simulation state to instead collect data through dynamic, partial, and potentially best-effort sampling akin to approaches traditionally used to study real-world systems (e.g., ice core samples, paleontological fossils). The aim of this strategy is to resolve scaling bottlenecks by economizing use of interconnect bandwidth, memory, and disk storage storage and better tolerate intermittent disruption. Trading a controlled amount of data detail for increased scalability and hardware accelerator compatibility would be highly worthwhile.
@@
Recently-developed hereditary stratigraphy methodology aims to bridge this gap by providing means for extracting phylogenetic information from distributed simulations that are efficient, robust, and straightforward to use (moreno2022hereditary). As is often the case in digital evolution, natural systems provide inspiration for the core strategy applied in hereditary stratigraphy: inference-based reconstruction. Natural history of biological life operates under no extrinsic provision for interpretable record-keeping, yet phylogenetic analysis of biological organisms has proved immensely fruitful. Such phylogenetic analyses are possible in biology because mutational drift encodes ancestry information in DNA genomes. Hereditary stratigraphy methods for decentralized work operate analogously, with ancestry information captured within agent genomes rather than through external tracking. The idea is to bundle agent genomes with special hereditary stratigraphic annotations in a manner akin to non-coding DNA (entirely neutral with respect to agent traits and fitness) and then use these annotations to perform phylogenetic reconstruction. The crux of hereditary stratigraphy algorithms, introduced in detail further on, is organization of genetic material to maximize reconstruction quality from a minimal memory footprint (moreno2022hereditary).
Since it was proposed, experimental work using hereditary stratigraphy has demonstrated viability in extracting information about underlying evolutionary conditions (moreno2024ecology), even at population scales reaching millions of agents/millions of generations using the 850,000 core Cerebras Wafer-Scale Engine hardware accelerator (moreno2024trackable). A number of options exist in configuring hereditary stratigraphy algorithms, but no work has yet systematically investigated how they relate to quality of phylogenetic reconstruction. In particular, it remains to be established how best to configure hereditary stratigraphy methodology to support use cases varying in scale, memory availability for annotation, and underlying evolutionary conditions. In this work, we report annotate-and-reconstruct experiments that evaluate reconstruction quality under possible hereditary stratigraphy configurations across a variety of use cases. We synthesize results from these experiments to suggest a prescriptive system of best practices for work with hereditary stratigraphy. Analysis covers three major configurable aspects of hereditary stratigraphy: (1) data structure implementation, (2) temporal data retention policy, and (3) size of stochastic lineage fingerprints. This work, in conjunction with availability of open-source software library utilities for hereditary stratigraphy (moreno2022hstrat), is hoped to catalyze means for phylogenetic analysis across a range of large-scale digital evolution projects.
@@ Methods
The goal of this work is to develop empirically informed best practice recommendations for using hereditary stratigraphy methods to trace ancestry relationships in digital evolution. This section begins with the conceptual basis of the hereditary stratigraphy approach, covering the checkpoint-based strategy used to measure relatedness. Guidance developed in this work, in particular, investigates two technical aspects of hereditary stratigraphy annotation:
@@
    checkpoint retention policy (steady vs. tilted), and
@@
    checkpoint storage strategy (column vs. surface).
@@
Having introduced the algorithmic structure of hereditary stratigraphy, attention turns to experiments conducted to evaluate which algorithmic configurations achieve best reconstruction quality. To be able to make general recommendations, tailored where necessary to application characteristics, we evaluated reconstruction quality across a variety of use case scenarios. Discussion covers the model used to generate reference phylogenies and the set of treatments evolutionary scenarios varying in scale and ground-truth phylogenetic richness explored. Finally, we describe metrics used to measure reconstruction quality, statistical methods, and software used for reconstruction quality experiments.
@@
Hereditary stratigraphy obtains phylogenetic information in a manner akin to how molecular phylogenetics approaches infer relatedness from comparisons among the genomes of biological organisms, which relies on the tendency for organisms with close hereditary relatedness to exhibit greater sequence similarity (yang2012molecular). However, phylogenetic analyses of sequence similarity under a mutational drift model is a nontrivial statistical problem (neyman1971molecular) that can be computationally demanding (konno2022deep; stamatakis2013novel), data-intensive (upwards of thousands of base pairs per genome) (parks2009increasing; cloutier2019whole; wortley2005much), and hindered by challenges arising from back mutation, mutational saturation, selection effects, and branch length differentials (brocchieri2001phylogenetic; moreira2000molecular). Although genome sites under mutational drift can be used to estimate relatedness among digital organisms (moreno2021case), a more lightweight and robust approach is desirable. To meet this need, hereditary stratigraphy systematizes a regimen of structured mutation that enables high-quality inference from small amounts of genetic material (moreno2022hereditary). The result is a general-purpose framework for phenotypically neutral “annotations” that can be affixed to digital organisms’ genomes, or individual genes, to make their lineages traceable (moreno2022hstrat).
@@
Hereditary stratigraphy works through a simple checkpointing procedure. In each generation, annotations are extended by appending a new random “fingerprint” value. These fingerprints, referred to as “differentiae” in the context of hereditary stratigraphy, serve to chronicle lineage history. The first pair of mismatching differentiae between two records definitively indicates a split in ancestry. Conversely, insofar as two annotations share identical differentiae, they likely share common ancestry. Under this framing, phylogeny reconstruction can be performed agglomeratively by successively percolating leaf taxa along the tree path of internal nodes consistent with their fingerprint sequence, then affixing them where common ancestry ends (moreno2024analysis).
@@
Because differentiae are randomly generated, it is possible that they collide by chance. The frequency of such spurious collisions depends on the size of fingerprint values used, e.g., a single bit, a byte, a 32-bit word, etc. Smaller differentiae reduce annotations’ memory footprint but make overestimation of relatedness more likely. Annotation memory use can also be reduced by discarding old differentiae as generations elapse. However, sparse retention reduces the number of reference points where divergence between lineages can be tested for. These two mechanisms enable tunable trade-offs between annotation size, inference precision, and inference accuracy. As such, strategies for differentia sizing and differentia retention are tested in reconstruction quality experiments, described below, to synthesize recommendations for best practice. The following section delves into the primary dimension of retention policy, steady versus tilted distribution, considered in this work.
@@
Figure 1: Steady versus tilted retention policy. Steady policy (top) retains differentia with time points spaced evenly across history. Tilted policy (bottom) retains differentia more densely over recent history, giving gap size proportional to time ago. Retained differentia are shown as filled diamonds and discarded differentia are shown as empty. Hybrid policy (not shown) allocates half of available space to hold tilted data and half to hold steady.
When pruning differentiae, care must be taken to ensure retention of checkpoint generations that maximize coverage across evolutionary history. In one possible strategy, retained time points would be spread evenly across history. We term this strategy as “steady” (han2005stream; zhao2005generalized). Such an approach ensures that last common ancestor (LCA) events can be discerned with consistent precision, no matter when they occurred.
Although the steady approach minimizes worst-case imprecision, there is reason to believe it may fail to allocate precision where it is most useful to discern phylogenetic topology. In most evolutionary scenarios, there is a consistent tendency for phylogenetic events concerning extant taxa to be more tightly packed in the near past (zhaxybayeva2004cladogenesis). A strategy accounting for this fact would retain newer time points at higher density than older time points. We call such an approach, where differentiae are retained in a recency-proportional manner, as “tilted” (han2005stream; zhao2005generalized). Preliminary experiments have indicated that, at least in some scenarios, annotations using tilted retention can yield higher quality phylogenetic reconstructions than with steady retention (moreno2022hereditary). Experiments in this work address this question more thoroughly, to assess the relative performance over a variety of evolutionary scenarios, including those expected to maintain greater amounts of ancient history. In addition, to assess whether benefits of these two approaches can be combined, some experiments consider a third, “hybrid”’ strategy, where half of annotation space is split evenly between steady and tilted strategies.
Figure 1 provides a visual comparison of steady and tilted retention strategies. The steady history spaces retained differentia at regular intervals, while gap sizes decrease with recency under the tilted policy. To fully appreciate the mechanics of differentia retention, however, an additional dimension of time must be considered. Rather than deciding retention just at one particular fixed point in time, these policies must handle continual accrual of new differentia as generations elapse. At each step, a new differentia is appended, and — as required — old differentia are discarded. Throughout, constraints on time point coverage and annotation size must be respected. Differentia curation for hereditary stratigraphy, indeed, turns out to be an instance of a larger class of “stream curation” problems where temporally representative records must be maintained on a rolling basis. For elaboration on this point and more on the implementation and algorithmic properties of retention policies, see moreno2024algorithms.
@@
Having just described approaches to deciding which differentiae should be stored, we now pivot to consider approaches to actually organize and store them. A suitable annotation data structure for differentiae curation should:
@@
    be readily serialized (to exchange annotated genomes between simulation processes), and
@@
The last point, minimizing representational overhead, is particularly critical given that typical use calls for single-bit and single-byte differentiae. Were 32- or 64-bit pointers or time point values required per differentia, bookkeeping overhead would greatly outweigh — and potentially crowd out — actual storage of lineage history information. As such, both approaches considered here — “column” and “surface”-based storage — pack differentiae in an array format and rely on positional context to identify them. Note that for such data to be readily legible, retention policies’ curated time points must be directly enumerable a priori for any arbitrary generation.
@@
The “column”’ approach arranges differentia in chronological order, with newest differentia stored last. This approach suits use of a dynamic array data structure (e.g., Python list/C++ std::vector) to store differentiae, as new additions can be accessioned through an append (e.g., “push_back”) operation. Accordingly, arbitrary curated collection growth can be supported. This allows for retention policies that provide hard guarantees for inference precision.  [^1] Hard fixed or recency-proportional bounds on differentia gap sizes require orders of growth in retention that are linear and logarithmic, respectively (moreno2024algorithms). One disadvantage to this approach, though, is that discarding old differentia requires a shift-down operation on all subsequent elements.
@@
The “surface” approach, in contrast, organizes differentiae directly onto a fixed-length buffer. Rather than appending as a “push_back” on the array, incoming differentiae are assigned an arbitrary buffer position and directly written there. One advantage of this approach is that separate garbage collection operations streamline away: new data simply overwrites that to be discarded. Another advantage is full use of available space: after the surface buffer is filled, it is guaranteed that stored differentiae fully utilize available capacity. Owing to discrepancy between projected upper bounds on retained size and actual usage, this is not the case for size-capped tilted retention using columns. However, to the surface’s disadvantage, dropping a level of abstraction to operate over buffer sites rather than differentia time points results in less fine-grained control over retention and, particularly in the case of size-capped steady retention, a somewhat looser adherence to idealized retention patterns. Additionally, by design, orders of growth beyond the surface’s fixed buffer size (e.g., logarithmic, linear) are not supported. For a more detailed description of motivation and implementation of surface-based algorithms, see (moreno2024trackable).
@@
In sum, the question of column- versus surface-based algorithms can be characterized as a trade-off between efficiency and exactitude. Indeed, we have found surface-based algorithms to provide order-of-magnitude speedups, as well as good compatibility with low-level, resource-constrained programming environments (notably, the Cerebras Wafer-Scale Engine hardware accelerator) (moreno2024trackable). To assess how, if at all, this trade-off impacts reconstruction quality, we include trials using both approaches in empirical annotate-and-reconstruct experiments, described below. Note that, in these experiments, we consider only fixed-size annotations. However, we anticipate nearly all hereditary stratigraphy use cases will apply fixed-size annotation, due to benefits from avoiding dynamic memory allocation and variable-length inter-process messaging at runtime.
@@
This section describes the evolution simulations used to generate reference phylogenies for empirical annotate-and-reconstruct experiments to appraise hereditary stratigraphy methods. To support the large, exact-tracked phylogenies needed for this purpose, we utilized a barebones evolution model. Genomes comprised a single floating-point value, with higher magnitude corresponding to higher fitness. Selection was performed using tournament selection with synchronous generations. Mutation was applied after selection, with a value drawn from a unit Gaussian distribution added to all genomes. To ensure asexual lineages, no crossover or recombination operations were performed. (Extensions of hereditary stratigraphy to sexual lineages are possible (moreno2024methods), but not explored in this work.) Evolutionary runs lasted 100,000 generations.
A major goal in these experiments is to assess reconstruction quality over a breadth of use case scenarios. To this end, we applied strong, explicit manipulations of evolutionary conditions to explore hereditary stratigraphy methods across regimes of phylogenetic structure. One focus was phylogenetic richness, the amount of distinct lineage history maintained within an extant population — also known as phylogenetic diversity (tucker2017guide). For a fixed population size, phylogenetic richness is increased by maintaining coexistence of deep phylogenetic branches. In two treatments, we applied spatial and ecological structure in concert to enhance phylogenetic richness (moreno2024ecology; gomez2019understanding; valiente2007facilitation), with spatial structure driven by a simple island population model and ecological structure driven by a simple niche model, respectively.
@@
The niche model, used to induce ecological structure, also applies a simple approach. Organisms were arbitrarily assigned to a niche at simulation startup, with fixed, equally-portioned population slots assigned to each niche. In the selection procedure, individuals exclusively compete with members of their own niche. Every generation, individuals swapped niches with probability $3.0517578125\times 10^{-8}$ (chosen so one niche swap would be expected every 500 generations at the larger population size and 4,000 generations at the smaller).
We also included a drift treatment, where selection was performed in a fully neutral manner. Drift conditions also enhance phylogenetic richness, but complements other surveyed treatments in operating through an entirely alternate mechanism.
Another objective of employing a highly-abstracted model to generate reference phylogenies is to enhance generality, in support of developing recommendations broadly applicable across digital evolution systems. Other work with this simple model supports its generality, finding generally analogous (albeit less accentuated) effects of manipulating evolutionary conditions on phylogenetic structure (moreno2024ecology). This model system has been established in other existing work, as well (moreno2023toward).
In addition to structural aspects of phylogeny composition, we also sought to understand the relationship between population scale and reconstruction quality. Population sizes of both 4,096 ($2^{12}$) and 65,536 ($2^{16}$) were used in experiments. Downsampling comprises an orthogonal dimension of phylogeny scale. We anticipate that most use cases of hereditary stratigraphy will collect and analyze only a subset of extant taxa for tractability of data collection, phylogenetic reconstruction, and phylogenetic analysis. As such, experiments downsampled generated annotations to 500 taxa. In experiments using larger population size, we also tested reconstructions over larger samples of 8,000 taxa.
@@
  • rich structure: tournament size 2 with 8 niches and 64 islands,
  • drift: tournament size 1 with no niching and no islands,
For each evolutionary regime, we tested five arrangements of annotation capacity and differentia size:
@@
Treatments also considered steady-versus-tilted retention policy and column-versus-surface implementation. Recall that these considerations are largely orthogonal, with the former determining the overall balance of recent-versus-ancient retention and the latter determining the particulars of discard sequencing. Experiments using surface implementation, in addition to steady and tilted approaches, additionally considered hybrid retention policy where buffer space was split evenly between the former approaches.
@@
To assess how hereditary stratigraphy would reconstruct surveyed reference trees, we simulated the inheritance of hereditary stratigraphic annotations along each reference phylogeny. This yielded a set of annotations equivalent to what would be attached to extant population members at the end of a run. Then, we used the agglomerative tree building implementation provided as hstrat.build_tree in the hstrat Python package (moreno2022hstrat). Thus, each reconstruction replicate has a directly corresponding reference tree from a perfect-tree treatment replicate.
Reconstruction was fast, taking less than a second for the 500-leaf tree with the 256-bit annotation and, in optimized mode, about 5 seconds to build the 8,000-leaf tree. We then a postprocessing step, hstrat.PeelBackConjoinedLeavesTriePostprocessor which accounts for the fact that entirely identical annotations may only arise due to spurious differentia collision because distinct leaf taxa by definition cannot have shared ancestry in their penultimate generation. This took about 20 seconds for the larger tree. Additional postprocessing options, including methods to assign timestamps to reconstructed inner nodes, are documented with the library.
@@
Assessment of reconstruction quality sought — in addition to characterizing a headline error measure across hereditary stratigraphy strategies — to provide diagnostic insight into the nature of reconstruction error produced and the underlying mechanistic reasons it occurs. Figure 2 depicts two distinct failure modes of hereditary stratigraphy-based reconstruction, which we would like to be able to distinguish. This example comprises three taxa: $A$, $B$, and $C$. In ground truth, $A$ and $B$ are most closely related and $C$ is an outgroup. We distinguish three classes of reconstruction outcomes:
@@
    correct reconstruction, where retained differentia suffice to distinguish the $(AB,C)$ branch then the subsequent $(A,B)$ branch;
@@
    incorrect reconstruction, where spurious differentia collision makes branches appear more closely related than they actually are, e.g., $B$ and $C$ sharing differentiae values by chance resulting in reconstruction where $B$ and $C$ are inferred as most closely related; and
@@
    unresolved reconstruction, where differentia necessary to distinguish branching order (i.e., between $(AB,C)$ and $(A,B)$) are not available but subsequent differentia collision does not occur, resulting in artifactual $(A,B,C)$ polytomy.

### v2

@@ Introduction
Ongoing advances in computing hardware have potential to unleash transformative, orders-of-magnitude growth in the scale and sophistication of agent-based evolutionary modeling and application-oriented evolutionary computation. For instance, emerging AI/ML accelerator hardware platforms (e.g., Cerebras Wafer-Scale Engine, Graphcore Intelligence Processing Unit, etc.) currently afford up to hundreds of thousands of processing cores within a single device (lauterbach2021path; jia2019dissecting). While these hardware platforms bear limitations characteristic of highly distributed, many-processor computation, their architecture is well-suited for work with agent-based models (ABM), as physical constraints in the layout of processor cores mirror the locally structured interactions typical in ABM. Nevertheless, significant challenges must be addressed to effectively harness highly distributed, many-processor computation for ABM workloads.
To advance on this front, we propose a fundamental re-frame of simulation that shifts from a paradigm of “complete,” deterministic observability of simulation state to instead collect data through dynamic, partial, and potentially best-effort sampling akin to approaches traditionally used to study real-world systems (e.g., ice core samples, paleontological fossils). The aim of this strategy is to resolve scaling bottlenecks by economizing use of interconnect bandwidth, memory, and disk storage storage and better tolerating intermittent disruption. Trading a controlled amount of data detail for increased scalability and hardware accelerator compatibility would be highly worthwhile.
@@
Recently developed “hereditary stratigraphy” methodology aims to bridge this gap by providing means for extracting phylogenetic information from distributed simulations that are efficient, robust, and straightforward to use (moreno2022hereditary). Natural history of biological life operates with no extrinsic provision for interpretable record-keeping, yet phylogenetic analysis of biological organisms has proved immensely fruitful. Such phylogenetic analyses are possible in biology because mutational drift encodes ancestry information in DNA genomes. Hereditary stratigraphy methods for decentralized work operate analogously, with ancestry information captured within agent genomes rather than through external tracking. The idea is to bundle agent genomes with special hereditary stratigraphic annotations in a manner akin to non-coding DNA (entirely neutral with respect to agent traits and fitness) and then use these annotations to perform phylogenetic reconstruction. The crux of hereditary stratigraphy algorithms, introduced in detail further on, is organization of genetic material to maximize reconstruction quality from a minimal memory footprint (moreno2022hereditary).
Since it was proposed, experimental work using hereditary stratigraphy has demonstrated viability in extracting information about underlying evolutionary conditions (moreno2024ecology), even at population scales reaching millions of agents/millions of generations using the 850,000 core Cerebras Wafer-Scale Engine hardware accelerator (moreno2024trackable). A number of options exist in configuring hereditary stratigraphy algorithms, but no work has yet systematically investigated how they relate to quality of phylogenetic reconstruction. In particular, it remains to be established how best to configure hereditary stratigraphy methodology to support use cases varying in scale, memory availability for annotation, and underlying evolutionary conditions. In this work, we report annotate-and-reconstruct experiments that evaluate reconstruction quality under possible hereditary stratigraphy configurations across a variety of use cases. We synthesize results from these experiments to suggest a prescriptive system of best practices for work with hereditary stratigraphy. Analysis covers three primary configurable aspects of hereditary stratigraphy: (1) data structure implementation, (2) temporal data retention policy, and (3) size of stochastic lineage fingerprints. This work, in conjunction with availability of open-source software library utilities for hereditary stratigraphy (moreno2022hstrat), is hoped to catalyze means for phylogenetic analysis across a range of large-scale digital evolution projects.
@@ Methods
The goal of this work is to develop empirically informed best practice recommendations for using hereditary stratigraphy methods to trace ancestry relationships in digital evolution. This section begins with the conceptual basis of the hereditary stratigraphy approach, covering the checkpoint-based strategy used to assess relatedness. Two technical aspects of hereditary stratigraphy annotation involved in this work are then introduced in corresponding subsections:
@@
    checkpoint retention policy (steady vs. tilted, Section Steady and Tilted Retention Algorithms), and
@@
    checkpoint storage strategy (column vs. surface, Section Column and Surface-Based Algorithms).
@@
Attention next turns to experiments conducted to evaluate the reconstruction quality obtained under alternate algorithmic configurations. For generality, we evaluated reconstruction quality across varied evolutionary scenarios. Discussion covers the model used to generate reference phylogenies and the set of treatments explored. Finally, we describe metrics used to measure reconstruction quality, statistical methods, and software used for reconstruction quality experiments.
@@
Hereditary stratigraphy delivers phylogenetic information analogously to molecular phylogenetics approaches, inferring lineage histories via the tendency for organisms with close hereditary relatedness to exhibit greater sequence similarity (yang2012molecular). However, phylogenetic analysis of sequence similarity under a mutational drift model is a nontrivial statistical problem (neyman1971molecular) that can be computationally demanding (konno2022deep; stamatakis2013novel), data-intensive (e.g., upwards of thousands of base pairs per genome) (parks2009increasing; cloutier2019whole; wortley2005much), and hindered by challenges arising from back mutation, mutational saturation, selection effects, and branch length differentials (brocchieri2001phylogenetic; moreira2000molecular). Although in digital evolution it is possible to use the coding content of model organisms’ genomes to estimate relatedness (moreno2021case), a more general and robust approach is desirable. To meet this need, hereditary stratigraphy defines a regimen of structured mutation that enables high-quality inference from small amounts of genetic material (moreno2022hereditary). The result is a general-purpose framework for phenotypically neutral “annotations” that can be affixed to digital organisms’ genomes, or even to individual genes, to make their lineages traceable (moreno2022hstrat).
@@
Hereditary stratigraphy composes annotations as a chronological sequence of checkpoint values. In each generation, annotations are extended by appending a new random “fingerprint” value. These fingerprints, referred to as “differentiae” in the context of hereditary stratigraphy, encode lineage history. The first pair of mismatching differentiae between two records definitively indicates a split in ancestry. Conversely, insofar as two annotations share identical differentiae, they likely share common ancestry. Under this framing, phylogeny reconstruction can be performed agglomeratively using a trie-based procedure. This approach successively percolates leaf taxa along the tree path of internal nodes consistent with their fingerprint sequence, affixing them where common ancestry ends (moreno2024analysis).
@@
Because differentiae are randomly generated, it is possible for annotations on diverged lineages to match by chance. The frequency of such spurious collisions depends on the size of fingerprint values used, e.g., a single bit, a byte, a 32-bit word, etc. Smaller differentiae reduce annotations’ memory footprint but make overestimation of relatedness more likely. Annotation memory use can also be reduced by discarding old differentiae as generations elapse. However, sparse retention reduces the number of reference points where divergence between lineages can be detected. These two mechanisms enable tunable trade-offs between annotation size, inference precision, and inference accuracy. As such, strategies for differentia sizing and differentia retention are tested in reconstruction quality experiments, described below, to synthesize recommendations for best practice. The following section delves into the primary dimension of retention policy, steady versus tilted distribution, considered in this work.
@@
Figure 1: Steady versus tilted retention policy. Steady policy (top) retains differentiae with time points spaced evenly across history. Tilted policy (bottom) retains differentiae more densely over recent history, giving gap size proportional to time ago. Retained differentia are shown as filled diamonds and discarded differentia are shown as empty. Hybrid policy (not shown) allocates half of available space to hold tilted data and half to hold steady.
When pruning differentiae, care must be taken to ensure retention of checkpoint generations that maximize coverage across evolutionary history. In one possible strategy, retained time points would be spread evenly across history. We term this strategy “steady” (han2005stream; zhao2005generalized). Such an approach ensures that last common ancestor (LCA) events can be discerned with consistent absolute precision, no matter when they occurred.
Although the steady approach minimizes worst-case imprecision, there is reason to believe it may fail to allocate precision where it would be most useful for discerning phylogenetic history. In most evolutionary scenarios, there is a general tendency for phylogenetic events concerning extant taxa to have occurred relatively recently (zhaxybayeva2004cladogenesis). A strategy accounting for this fact would retain newer time points at higher density than older time points. We refer to such an approach, where differentiae are retained in a recency-proportional manner, as “tilted” (han2005stream; zhao2005generalized).
Preliminary experiments have indicated that, in some scenarios, annotations using tilted retention can yield higher quality phylogenetic reconstructions than those using steady retention (moreno2022hereditary). Experiments here address this question more thoroughly; we assess the relative performance over a variety of evolutionary scenarios, including those expected to maintain greater amounts of ancient history. In addition, to assess whether benefits of these two approaches can be combined, some experiments consider a third, “hybrid” strategy, where annotations are split half-and-half between steady and tilted strategies.
Figure 1 provides a visual comparison of steady and tilted retention strategies. The steady strategy retains differentia at regular intervals, while the tilted policy retains differentia in proportion to their recency. Note that these policies must handle continual accrual of new differentia as generations elapse. At each step, a new differentia is appended, and — as required — old differentia are discarded. Throughout, constraints on time point coverage and annotation size must be respected. For more detail on algorithmic aspects of this process, see moreno2024algorithms; moreno2024structured.
@@
Having just described approaches to deciding which differentiae should be stored, we next consider implementation-level approaches to organize and store them in practice. A suitable annotation data structure for differentiae curation should:
@@
    be readily serialized (to exchange annotated genomes between processors), and
@@
The last point, minimizing representational overhead, is particularly critical given that typical use calls for single-bit and single-byte differentiae. Requiring 32- or 64-bit pointers or time point values per differentia would cause bookkeeping overhead to greatly outweigh — and potentially crowd out — useful lineage history information. As such, both approaches considered here — “column” and “surface”-based storage — pack differentiae in an array format and rely on positional context to identify them. For such data to be readily legible, retention policies’ curated time points must be directly enumerable a priori for any arbitrary generation.
@@
The “column” approach arranges differentiae in chronological order, with newest differentiae stored last. This approach suits use of a dynamic array data structure (e.g., Python list/C++ std::vector), as new additions can be accessioned through an append (e.g., “push_back”) operation. Accordingly, the “column” approach supports arbitrary growth in curated collection size. This property allows for retention policies that provide hard guarantees on inference precision.  [^1] Hard fixed or recency-proportional bounds on differentia gap sizes require orders of growth in retention that are linear and logarithmic, respectively (moreno2024algorithms). One disadvantage to this approach, though, is that discarding old differentia requires an $\mathcal{O}(n)$ shift-down operation on all subsequent elements.
@@
The “surface” approach, in contrast, organizes differentiae directly onto a fixed-length buffer (moreno2024structured). Rather than appending as a “push_back” on the array, incoming differentiae are assigned an arbitrary buffer position and directly written there in $\mathcal{O}(1)$ time. One advantage of this approach is that explicit garbage collection operations are unnecessary; new data simply overwrites that to be discarded. Another advantage is full use of available space — after the surface buffer is filled, it is guaranteed that stored differentiae fully utilize available capacity. Owing to discrepancy between projected upper bounds on retained size and actual usage, this guarantee does not hold for column-based tilted retention where maximum size is capped. However, to the surface approach’s disadvantage, dropping a level of abstraction to operate over buffer sites rather than differentia time points results in less fine-grained control over retention and a somewhat looser adherence to idealized retention patterns. Additionally, by design, growth beyond the surface’s fixed buffer size is not supported. For a more detailed description of motivation for surface-based algorithms, see (moreno2024trackable).
@@
In sum, the question of column- versus surface-based algorithms can be characterized as a trade-off between efficiency and exactitude. Indeed, surface-based algorithms provide order-of-magnitude speedups, as well as better compatibility with low-level, resource-constrained programming environments (moreno2024trackable). To assess how, if at all, this trade-off impacts reconstruction quality, we include trials using both approaches in empirical annotate-and-reconstruct experiments, described below. Note that, in these experiments, we consider only fixed-size annotations. However, we anticipate nearly all hereditary stratigraphy use cases will apply fixed-size annotation due to benefits of avoiding dynamic memory allocation and variable-length inter-process messaging at runtime.
@@
This section describes the evolution simulations used to generate reference phylogenies for empirical annotate-and-reconstruct experiments. To support the large, exact-tracked phylogenies needed for this purpose, we used a simple evolution model. Genomes comprised a single floating-point value, with higher magnitude corresponding to higher fitness. We used tournament selection with synchronous generations. Mutation was applied after selection, with a value drawn from a unit Gaussian distribution added to all genomes. Experiments used asexual reproduction, with no crossover or recombination. (Extensions of hereditary stratigraphy to sexual lineages are possible (moreno2024methods), but not explored in this work.) Evolutionary runs lasted 100,000 generations.
A key consideration in our experiments is assessing reconstruction quality over a breadth of evolutionary scenarios. To this end, we applied strong, explicit manipulations of evolutionary conditions to explore hereditary stratigraphy methods across diverse regimes of phylogenetic structure. One focal variable was phylogenetic diversity, the amount of distinct lineage history maintained within an extant population (tucker2017guide). For a fixed population size, phylogenetic richness is increased by coexistence of deep phylogenetic branches. We explored treatments that promote phylogenetic richness via spatial and ecological structure (moreno2024ecology; gomez2019understanding; valiente2007facilitation). Spatial structure in experiments was implemented as a simple island population model and ecological structure was implemented as a simple niche model, respectively.
@@
The niche model, used to induce ecological structure, also applies a simple approach. Organisms were arbitrarily assigned to a niche at simulation startup, with fixed, equally-portioned population slots assigned to each niche. In the selection procedure, individuals exclusively compete with members of their own niche. Every generation, individuals swapped niches with probability $3.05\times 10^{-8}$ (chosen so one niche swap would be expected every 500 generations at the larger population size and 4,000 generations at the smaller).
We also included a drift treatment, where selection was performed in a fully neutral manner. Drift conditions also enhance phylogenetic richness, but complement other surveyed treatments in operating through a separate mechanism.
Another objective of employing a highly abstracted model to generate reference phylogenies is generality. Other work with this simple model suggests that effects on phylogenetic structure induced under treatments surveyed here generalize across model systems, albeit to a less accentuated degree (moreno2024ecology). This model system has been established in other existing work, as well (moreno2023toward).
In addition to structural aspects of phylogeny composition, we also sought to understand the relationship between population scale and reconstruction quality. Population sizes of both 4,096 ($2^{12}$) and 65,536 ($2^{16}$) were used in experiments. We anticipate that many use cases of hereditary stratigraphy will collect and analyze only a subset of extant taxa for tractability of data collection, phylogenetic reconstruction, and phylogenetic analysis. As such, experiments downsampled generated annotations to 500 taxa. In experiments using larger population size, we also tested reconstructions over larger samples of 8,000 taxa.
@@
  • rich structure: tournament size 2 with 8 niches and 64 islands, and
  • drift: tournament size 1 with no niching and no islands.
For each evolutionary regime, we tested four arrangements of annotation capacity and differentia size:
@@
Treatments also considered steady-versus-tilted retention policy and column-versus-surface implementation. Recall that the former determines the overall balance of recent-versus-ancient checkpoints kept and the latter determines the implementation-level particulars of discard sequencing. Experiments using surface implementation, in addition to steady and tilted approaches, also considered a steady/tilted hybrid retention policy.
@@
To assess how hereditary stratigraphy would reconstruct ground-truth reference trees, we simulated the inheritance of hereditary stratigraphic annotations along reference phylogenies. This yielded a set of annotations equivalent to what would be attached to extant population members at the end of a run. Then, we used the agglomerative tree building implementation provided as hstrat.build_tree in the hstrat Python package (moreno2022hstrat). Thus, each reconstruction replicate has a directly corresponding reference tree from an exactly tracked evolution run.
Reconstruction was fast, taking less than a second for the 500-leaf tree with the 256-bit annotation and, in optimized mode, about 5 seconds to build the 8,000-leaf tree. We then apply a postprocessing step, hstrat.PeelBackConjoinedLeavesTriePostprocessor which accounts for the fact that entirely identical annotations may only arise due to spurious differentia collision because distinct leaf taxa by definition cannot have shared ancestry in their penultimate generation. This took about 20 seconds for the larger tree. Additional postprocessing options, including methods to assign timestamps to reconstructed inner nodes, are documented with the library.
@@
Assessment of reconstruction quality sought — in addition to characterizing an overall error measure across hereditary stratigraphy strategies — to provide diagnostic insight into the nature of reconstruction error produced and the underlying mechanistic reasons it occurs. Figure 2 depicts two distinct failure modes of hereditary stratigraphy-based reconstruction, which we seek to distinguish. This example involves three taxa: $A$, $B$, and $C$. In ground truth, $A$ and $B$ are most closely related and $C$ is an outgroup. We distinguish three classes of reconstruction outcomes:
@@
    correct reconstruction, where retained differentiae suffice to distinguish the $(AB,C)$ branch then the subsequent $(A,B)$ branch;
@@
    incorrect reconstruction, where spurious differentia collision makes branches appear more closely related than they actually are, e.g., $B$ and $C$ sharing differentiae values by chance, resulting in reconstruction where $B$ and $C$ are inferred as most closely related; and
@@
    unresolved reconstruction, where differentia necessary to distinguish branching order are not available but subsequent differentia collision does not occur, resulting in artifactual $(A,B,C)$ polytomy.